ok, last one, honest! :>
(I think I'm getting too old for this!)
Two medium eggs and three cups of ice cream contain 701 mg of cholesterol. One med. egg and 1 c. ice cream contains 325 mg cholesterol. Find the cholesterol content of one egg and for one cup ice cream. Thanks!!

Answer by dipti1(1)

2 medium eggs + 3 cups Ice cream = 701 mg. Cholesterol
2 medium eggs + 2 cups Ice Cream = 650 mg. Cholesterol
_______________________________________________________
0 + 1 cup Ice Cream = (701 - 650) - 51 mg. Cholesterol
Now, substitute the value of 1 cup of Ice cream in the follwoing formula :
1 medium egg + 1 cup of Ice Cream = 325 mg. Cholesterol
so,
1 medium egg + 51 mg. cholesterol = 325 mg. Cholesterol
so 1 medium egg = 325 - 51 mg. Cholesterol

1 medium egg = 274 mg. Cholesterol
Thus 1 cup of Ice Cream = 51 mg. Cholesterol &
1 medium egg = 274 mg. Cholesterol